### v1

# Newton Methods for Mean Field Games: A Numerical Study

Abstract.
We address the numerical solution of second-order Mean Field Game problems through Newton iterations in infinite dimensions, introduced in [newton], where quadratic convergence of the method was rigorously established. Building upon this theoretical framework, we develop new numerical discretization techniques, including both a finite difference and a semi-Lagrangian scheme, that enable an effective computational implementation of the infinite-dimensional iterations. The proposed methods are tested on several benchmark problems, and the resulting numerical experiments demonstrate their robustness, accuracy, and efficiency. A comparative analysis between the two schemes and existing approaches from the literature is also presented, highlighting the potential of Newton-based solvers for MFG systems.
1Dipartimento di Matematica Guido Castelnuovo, Sapienza Università di Roma, 00185 Rome, Italy (carlini@mat.uniroma1.it).
2Faculty of Mathematics, University of Vienna, Oskar-Morgenstern-Platz 1, 1090 Vienna, Austria. (ahmad.zorkot@univie.ac.at).
1. Introduction
Mean field games (MFG) offer a powerful framework for modeling the strategic behavior of a large number of indistinguishable agents who interact through the statistical distribution of their states. Introduced independently by Lasry and Lions [ll2, ll3, L-L] and by Huang, Malhamé, and Caines [malhame2, malhame4], MFG has become a central tool for the analysis of Nash equilibria in stochastic differential games with infinitely many players. MFG models have found applications in a broad range of fields, including economics, finance, traffic flow, crowd dynamics, and the social sciences. We refer to [carmonafinance, ext6, mfgnash, mfgnash2, MR3195844] for general overviews of the theory and its applications.
From an analytical viewpoint, the classical formulation of MFGs introduced in [ll2, ll3] consists of a coupled system of partial differential equations (PDEs). The backward component is a Hamilton–Jacobi–Bellman (HJB) equation associated with the optimal control problem of a representative agent, complemented by a terminal condition, while the forward component is a Fokker–Planck (FP) equation governing the evolution of the population density and subject to an initial condition. Comprehensive presentations of this framework can be found in the monographs [ext7, ext6, mfgnash2], the survey [MR3195844], and the lecture notes [MR4214773].
In this work, we focus on the following time-dependent, second-order Mean Field Game system with local coupling, posed on the $d$-dimensional torus $\mathbb{T}^{d}$:
(1.1) $\begin{cases}-\partial_{t}u-\nu\Delta u+H(x,Du)=F(m(t,x))&\text{in }[0,T]\times\mathbb{T}^{d},\\ \partial_{t}m-\nu\Delta m-{\rm{div}}(mD_{p}H(x,Du))=0&\text{in }[0,T]\times\mathbb{T}^{d},\\ m(0,x)=m_{0}(x),u(T,x)=G(x)&\text{in }\mathbb{T}^{d},\end{cases}$
where $T>0$ denotes the time horizon, $\nu>0$ is the diffusion coefficient, $H$ is a convex Hamiltonian, and $F$ is a coupling term depending locally on the density $m$. The choice of the torus $\mathbb{T}^{d}=\mathbb{R}^{d}/\mathbb{Z}^{d}$ allows us to avoid the treatment of boundary conditions.
The numerical approximation of MFG systems has been the subject of intensive research in recent years. For second-order (stochastic) problems such as (1.1), finite difference schemes have been investigated in [finitedifference1, finitediff2, MR3452251], while semi-Lagrangian approaches have been developed in [CSorder2, Calzola_Carlini_Silva_lg_second_order]. More recently, data-driven techniques, including deep learning, reinforcement learning, and tensor-based methods, have been applied to MFG models in [MR4264647, MR4522347, MR4440805, carlinisaluzzi2025]. In the first-order (deterministic) setting, we refer to [CSorder1, carlini2023lagrangegalerkin, saeed, gianatti2023approximation, GS, carlinicoscetti2025]. Broader surveys of numerical methods for MFGs can be found in [mathieuachdou, mathieu1] and the references therein.
The main focus of the present work is the application of Newton’s method in infinite-dimensional function spaces to the continuous MFG system (1.1). This approach, originally proposed in [newton] (see also [berry-silva] for the stationary case), reformulates the coupled PDE system as a nonlinear operator equation and constructs successive approximations by linearization around the current iterate. Each Newton step therefore requires the solution of a coupled forward–backward linear parabolic system. A key theoretical result established in [newton] is the quadratic local convergence of the continuous Newton iterations under suitable regularity and proximity assumptions.
For clarity of presentation, we introduce the discretization in the case of a separable Hamiltonian. Nevertheless, the proposed methodology naturally extends to non-separable Hamiltonians, and this extension is illustrated through a dedicated numerical experiment.
To implement Newton’s method for system (1.1), we introduce the operator
(1.2) ${\mathcal{F}}:(u,m)\to\left(\begin{array}[]{l}-\partial_{t}u-\nu\Delta u+H(\cdot,Du)-F(m)\\ \partial_{t}m-\nu\Delta m-{\rm{div}}(mD_{p}H(\cdot,Du))\\ u(T,\cdot)-G(\cdot)\\ m(0,\cdot)-m_{0}(\cdot)\end{array}\right),$
so that system (1.1) can be written compactly as
$\mathcal{F}(u,m)=0.$
Assuming sufficient smoothness of the data, the Newton iterates $(u^{n},m^{n})$ for $n\geq 1$ are defined by
(1.3) $J{\mathcal{F}}(u^{n-1},m^{n-1})(u^{n}-u^{n-1},m^{n}-m^{n-1})=-{\mathcal{F}}(u^{n-1},m^{n-1}),$
where $J{\mathcal{F}}$ denotes the Jacobian of $\mathcal{F}$}, given explicitly by
(1.4) $J{\mathcal{F}}(u,m)(v,\rho)=\left(\begin{array}[]{lr}-\partial_{t}v-\nu\Delta v+D_{p}H(\cdot,Du)Dv\quad&-F^{\prime}(m)\rho\\[4.0pt] -{\rm{div}}(mD_{pp}H(\cdot,Du)Dv)&\partial_{t}\rho-\nu\Delta\rho-{\rm{div}}(D_{p}H(\cdot,Du)\rho)\\ v(T,\cdot)&0\\ 0&\rho(0,\cdot)\end{array}\right),$
for Hölder continuous functions $v$ and $\rho$ defined on $0,T]\times\mathbb{R}^{d}$.
Setting $q^{n}=D_{p}H(\cdot,Du^{n-1})$ and substituting (1.4) into (1.3), we obtain the linearized system
(1.5) $\begin{cases}-\partial_{t}u^{n}-\nu\Delta u^{n}+q^{n}Du^{n}=q^{n}Du^{n-1}-H(x,Du^{n-1})+F(m^{n-1})\\ \hskip 142.26378pt+F^{\prime}(m^{n-1})(m^{n}-m^{n-1})\quad&\text{in }[0,T]\times\mathbb{T}^{d},\\[6.0pt] \partial_{t}m^{n}-\nu\Delta m^{n}-{\rm{div}}(m^{n}q^{n})={\rm{div}}(m^{n-1}D_{pp}H(x,Du^{n-1})(Du^{n}-Du^{n-1}))\quad&\text{in }[0,T]\times\mathbb{T}^{d},\\[6.0pt] m^{n}(x,0)=m_{0}(x),\quad u^{n}(x,T)=G(x)\quad&\text{in }\mathbb{T}^{d}.\end{cases}$
It was shown in [newton] that, under appropriate assumptions, the sequence generated by these Newton iterations converges locally to the unique solution of (1.1), with a quadratic rate provided the initial guess is sufficiently close to the exact solution. Despite these strong theoretical guarantees, the numerical behavior of the infinite-dimensional Newton method has not been extensively investigated. The aim of the present work is to address this gap by designing suitable discretization schemes and assessing their performance through numerical experiments.
To discretize and solve the linearized systems (1.5), we consider two complementary numerical strategies: a semi-Lagrangian (SL) method and an implicit finite difference scheme. For both approaches, we establish well-posedness results and evaluate their efficiency and accuracy numerically. We also compare our methodology with the Newton-based finite difference scheme introduced in [finitedifference1], highlighting the distinction between the paradigms of “iterate then discretize”, where Newton’s method is applied at the continuous level, and “discretize then iterate”, where Newton’s method is applied directly to the discrete nonlinear system. Our experiments indicate that both strategies yield comparable accuracy and computational cost, while the former offers a simpler and more flexible discretization framework.
In contrast to approaches such as [finitedifference1, CSorder2], which discretize the fully nonlinear MFG system (1.1) and subsequently apply Newton’s method at the discrete level, our strategy applies standard finite difference and semi-Lagrangian discretizations only to the linear systems arising at each Newton iteration. This results in a significantly simpler numerical structure. Moreover, our results show that the semi-Lagrangian scheme exhibits greater robustness in the hyperbolic regime, corresponding to small values of $\nu$. Finite-difference-based schemes, on the other hand, may experience numerical instabilities in this setting, in agreement with known limitations discussed in [achdou1, achdou2], and often require continuation strategies to gradually decrease $\nu$.
Since Newton’s method is only locally convergent, its effectiveness strongly depends on the choice of the initial guess. To enhance robustness when standard Newton iterations fail, we incorporate a globalized Newton strategy based on a line-search procedure, following [global-newton]. This globalization technique, driven by a suitable merit function, proves particularly effective for the finite-difference-based Newton solver in the hyperbolic regime.
The remainder of the paper is organized as follows. Section 2 is devoted to the construction of a semi-Lagrangian discretization of the linearized system. In Section 3, we introduce an implicit finite difference scheme. Finally, Section 4 presents the numerical experiments and a comparative analysis of the proposed methods.
2. A semi-Lagrangian scheme
In this section, we discretize the iterative system (1.5) by means of a semi-Lagrangian scheme in the two dimensional state-space and we prove the well-posedness of the discrete system.
We refer to [Calzola_Carlini_Silva_lg_second_order, CSorder2] for the early work on approximating the second-order MFG systems using a semi-Lagrangian scheme, and to [CC] for a semi-Lagrangian scheme applied to parabolic equations.
2.1. Notations and definitions
In what follows, to simplify the discussion, we take $d=2$ and consider the Hamiltonian
(2.1) $H(x,p)=\frac{|p|^{2}}{2}-V(x),\qquad\text{for }(x,p)\in\mathbb{T}^{2}\times\mathbb{R}^{2}.$
where $V$ is a given bounded potential.
Given two positive integers $N_{t}$ and $N_{h}$, we define the time step $\Delta t=\frac{T}{N_{t}}$ and the space step $h=\frac{1}{N_{h}}$. We also introduce the index sets
$\mathcal{I}_{\Delta t}:=\{0,\dots,N_{t}\},\qquad\mathcal{I}_{\Delta t}^{*}:=\mathcal{I}_{\Delta t}\setminus\{N_{t}\},\qquad\mathcal{I}_{h}:=\{0,\dots,N_{h}-1\}.$
We define the discrete time grid $\mathcal{G}_{\Delta t}:=\{t_{k}=k\Delta t\mid k\in\mathcal{I}_{\Delta t}\}$ and discrete space grid $\mathcal{G}_{h}:=\{x_{i,j}=(ih,jh)\mid i,j\in\mathcal{I}_{h}\}$. We denote by $B(\mathcal{G}_{h})$ and $B(\mathcal{G}_{\Delta t}\times\mathcal{G}_{h})$ the two sets of functions defined in $\mathcal{G}_{h}$ and $\mathcal{G}_{\Delta t}\times\mathcal{G}_{h}$ respectively.
The objective is to construct two discrete functions $u^{n,k}_{[i,j]}$ and $m^{n,k}_{[i,j]}$ approximating the solution $(u^{n},m^{n})$ of (1.5) at the grid points $(t_{k},x_{i,j})$ for all $k\in\mathcal{I}_{\Delta t}$ and $i,j\in\mathcal{I}_{h}$. The index operator $\cdot$$\cdot$], defined by $i,j]=\big((i+N_{h})\bmod N_{h},\,(j+N_{h})\bmod N_{h}\big)$, accounts for periodic boundary conditions. For notational simplicity we will also write $(u^{n,k}_{i,j},m^{n,k}_{i,j})=(u^{n,k}_{[i,j]},m^{n,k}_{[i,j]})$. Given a grid function $v\in\mathcal{B}(\mathcal{G}_{h})$, we introduce the first order central differences operators
(2.2) $\begin{array}[]{rcl}\displaystyle(D_{1}v)_{i,j}&=\displaystyle\frac{v_{i+1,j}-v_{i-1,j}}{2h}\quad\text{for }i,j\in\mathcal{I}_{h},\\[10.0pt] \displaystyle(D_{2}v)_{i,j}&=\displaystyle\frac{v_{i,j+1}-v_{i,j-1}}{2h}\quad\text{for }i,j\in\mathcal{I}_{h},\end{array}$
and define the discrete gradient operator $D_{h}$ as
(2.3) $(D_{h}v)_{i,j}=((D_{1}v)_{i,j},(D_{2}v)_{i,j})^{\top}\quad\text{for }i,j\in\mathcal{I}_{h}.$
Then, for a discrete vector field $p=(p_{1},p_{2})\in B(\mathcal{G}_{h})^{2}$, we define the discrete divergence operator ${\rm{div}}_{h}$ as
(2.4) $({\rm{div}}_{h}(vp))_{i,j}=\frac{1}{2h}\Big(v_{i+1,j}(p_{1})_{i+1,j}-v_{i-1,j}+v_{i,j+1}(p_{2})_{i,j+1}-v_{i,j-1}(p_{2})_{i,j-1}\Big),\quad i,j\in\mathcal{I}_{h}.$
Given $\phi\in\mathcal{B}(\mathcal{G}_{h})$, we define its piecewise linear interpolant as
$I[\phi](x)=\sum_{i,j\in\mathcal{I}_{h}}\beta_{i,j}(x)\phi_{i,j}\quad\text{for all }x\in\mathbb{T}^{2},$
where $\phi_{i,j}=\phi(x_{i,j})$, and $\beta_{i,j}$ denote the piecewise linear finite element basis associated to the grid $\mathcal{G}_{h}$. For every $\varphi\colon\mathbb{T}^{2}\to\mathbb{R}$ denote by $\varphi|_{\mathcal{G}_{h}}$ its restriction to $\mathcal{G}_{h}$. If $\varphi$ is of class $C^{2}$ and has bounded second order derivatives, it follows from [quarteronivalli94, Remark 3.4.2] that
(2.5) $\|\varphi(\cdot)-I[\varphi|_{\mathcal{G}_{h}}](\cdot)\|_{\infty}\leq Ch^{2},$
where $C>0$ depends only on $\varphi$.
2.2. Semi-Lagrangian scheme for the backward equation
Building on the notations and framework introduced in the previous section, we now apply the semi–Lagrangian discretization to approximate the iterative system (1.5). We start with the backward equation, which can be written as follows:
(2.6) $\begin{cases}-\partial_{t}u^{n}-\frac{\sigma^{2}}{2}\Delta u^{n}+q^{n}Du^{n}-G^{n}(t,x)=0\quad&\text{in }[0,T]\times\mathbb{T}^{2},\\ u^{n}(x,T)=G(x)\quad&\text{in }\mathbb{T}^{2},\end{cases}$
where
$\sigma=\sqrt{2\nu},\quad q^{n}(t,x)=Du^{n-1}(t,x)$
and
$\begin{array}[]{rcl}G^{n}(t,x)&=&\displaystyle\frac{|q^{n}(t,x)|^{2}}{2}-V(x)\\[6.0pt] &&+\displaystyle F(m^{n-1}(t,x))+F^{\prime}(m^{n-1}(t,x))(m^{n}(t,x)-m^{n-1}(t,x)),\quad\text{for }(t,x)\in[0,T]\times\mathbb{T}^{2}.\end{array}$
By the Feynman-Kac formula (see e.g [yong]), the solution $u^{n}$ to (2.6), admits the following representation
(2.7) $u^{n}(t,x)=\mathbb{E}\Big[\int_{t}^{T}G^{n}(s,X^{t,x}(s)){\rm d}s+G(X^{t,x}(T))\Big],\quad\text{for }(t,x)\in[0,T]\times\mathbb{T}^{2},$
where $X^{t,x}$ denotes characteristics solving
(2.8) $\left\{\begin{aligned} dX(s)&=q^{n}(s,X(s))+\sigma dW(s)\quad\text{for }s\in(t,T),\\ X(t)&=x.\end{aligned}\right$.
We explain now how to construct a SL approximation using the technique shown in [falcone]. First, notice that (2.7) imply that for every $k\in\mathcal{I}^{*}_{\Delta t}$, we have
(2.9) $u^{n}(t_{k},x)=\mathbb{E}\Big[\int_{t_{k}}^{t_{k+1}}G^{n}(s,X^{t_{k},x}(s)){\rm d}s+u^{n}(t_{k+1},X^{t_{k},x}(\Delta t))\Big].$
We denote by $q^{n,k}_{i,j}\in\mathbb{R}^{2}$ an appropriate approximation of the drift term $q^{n}(t_{k},x_{i,j})$ for each $k\in\mathcal{I}_{\Delta t}$ and $i,j\in\mathcal{I}_{h}$. Then, we approximate the expectation in (2.9) (see e.g [Kloeden]) as
(2.10) $\mathbb{E}\left[u^{n}(t_{k},X^{t_{k},x_{i,j}}(\Delta t))\right]=\frac{1}{4}\sum_{\ell=1}^{4}u^{n}(t_{k},y^{\ell}(x_{i,j},q^{n,k}_{i,j}))+O((\Delta t)^{2}),$
where
(2.11) $y^{\ell}(x_{i,j},q^{n,k}_{i,j})=(x_{i,j}+\Delta tq^{n,k}_{i,j}+\sqrt{2\Delta t}\sigma e^{\ell})_{p}\quad\text{for }\ell=1,\ldots,4,$
with $e^{\ell}$ representing four vectors of $\mathbb{R}^{2}$ with one component equal to $\pm 1$ and the other null, and
$(z)_{p}=(z_{1}-\lfloor z_{1}\rfloor,z_{2}-\lfloor z_{2}\rfloor)\quad\text{for all }z=(z_{1},z_{2})\in\mathbb{R}^{2},$
denotes the periodic projection on $\mathbb{T}^{2}$ of $z\in\mathbb{R}^{2}$.
Finally, by combining (2.7), (2.9), and (2.10), and using the rectangular rule to approximate the integral term in (2.9), we define the semi–Lagrangian scheme for equation (2.6) as follows. Given $q^{n}\in B(\mathcal{G}_{\Delta t}\times\mathcal{G}_{h})^{2}$ and $m^{n},m^{n-1}\in B(\mathcal{G}_{\Delta t}\times\mathcal{G}_{h})$, find $u^{n}\in B(\mathcal{G}_{\Delta t}\times\mathcal{G}_{h})$ such that
[…]
where, for every $f\in B(\mathcal{G}_{h})$, $k\in\mathcal{I}_{\Delta t}^{*}$, and $i,j\in\mathcal{I}_{h}$,
$\begin{split}S^{n}_{k,(i,j)}(f):=\frac{1}{4}\sum_{\ell=1}^{4}I[f](y^{\ell}(x_{i,j},q^{n,k}_{i,j}))+\Delta tG^{n}(t_{k},x_{i,j}).\end{split}$
Let $\phi\in C^{2}_{0}(\mathbb{T}^{2})$, and let us denote by $C$ a positive real number which can depend only on $\phi$. From (2.11) and Taylor expansion (see e.g [CC]) we have
(2.13) $\Big|\frac{1}{4}\sum_{\ell=1}^{4}\phi(y^{\ell}(x_{i,j},q^{n,k}_{i,j}))-\left(\phi(x_{i,j})+\Delta t\frac{\sigma^{2}}{2}\Delta\phi(x_{i,j})+\Delta tq^{n,k}_{i,j}D\phi(x_{i,j})\right)\Big|\leq C(\Delta t)^{2}\quad\text{for all }i,j\in\mathcal{I}_{h}.$
Let $u^{n}$ be a smooth solution to (2.6), with bounded derivatives. Then, using (2.5) and (2.13), for every $k\in\mathcal{I}_{\Delta t}^{*}$ and $i,j\in\mathcal{I}_{h}$, the consistency error of scheme (2.12) satisfies
$T_{\Delta t,h}(t_{k},x_{i,j}):=\frac{1}{\Delta t}\left(u^{n}(t_{k+1},x_{i,j})-S^{n}_{k,(i,j)}(u^{n}(t_{k+1})\right)=O\Big((\Delta t)^{2}+\frac{h^{2}}{\Delta t}\Big)\quad(t,x)\in(0,T)\times\mathbb{T}^{2}.$
2.3. A semi-Lagrangian scheme for the forward equation
Let us now consider the second equation in system (1.5)
(2.14) $\begin{cases}\partial_{t}m^{n}-\frac{\sigma^{2}}{2}\Delta m^{n}-{\rm{div}}(m^{n}q^{n})={\rm{div}}(m^{n-1}(t,x)(Du^{n}(t,x)-Du^{n-1}(t,x)))&\text{in }[0,T]\times\mathbb{T}^{2},\\ m^{n}(0,x)=m_{0}(x)&\text{in }\mathbb{T}^{2}.\end{cases}$
Following the same analogue in [CSorder2], we propose the following scheme to approximate (2.14). Given $q^{n},D_{h}u^{n-1},D_{h}u^{n}\in B(\mathcal{G}_{\Delta t}\times\mathcal{G}_{h})^{2}$, and $m^{n-1}\in B(\mathcal{G}_{\Delta t}\times\mathcal{G}_{h})$, find $m^{n}\in B(\mathcal{G}_{\Delta t}\times\mathcal{G}_{h})$ such that
(2.15) $\begin{cases}\begin{split}m^{n,k+1}_{i,j}&=(S^{n}_{k,(i,j)})^{*}(m^{n,k})\quad\text{for all }k\in\mathcal{I}^{*}_{\Delta t},\,i,j\in\mathcal{I}_{h},\\[4.0pt] m^{n,0}_{i,j}&=m_{0}(x_{i,j}),\end{split}\end{cases}$
where, for a given function $f\in B(\mathcal{G}_{h})$, $k\in\mathcal{I}_{\Delta t}^{*}$ and $i,j\in\mathcal{I}_{h}$
$\begin{split}(S^{n}_{k,(i,j)})^{*}(f)=\frac{1}{4}\sum_{\ell=1}^{4}I^{*}[f](y^{\ell}(x_{i,j},q^{n,k}_{i,j}))+\Delta t{({\rm{div}}_{h}(m^{n-1,k+1}({D}_{h}u^{n,k+1}-D_{h}u^{n-1,k+1})))_{i,j}},\end{split}$
where, for every $i,j\in\mathcal{I}_{h}$, $I^{*}[f](y^{\ell}(x_{i,j},q^{n,k}_{i,j}))$ is the adjoint operator of $f\to I[f](y^{\ell}(x_{i,j},q^{n,k}_{i,j}))$.
2.4. The discrete Newton system
Given $v\in\mathcal{B}(\mathcal{G}_{h})$, we denote by $V\in\mathbb{R}^{{N_{h}^{2}}}$ the vectors such that
(2.16) $V_{i+jN_{h}}=v_{i,j}\quad\forall\quad i,j\in\mathcal{I}_{h}.$
Then the discrete gradient operator $D_{h}:\mathbb{R}^{N_{h}}\to\mathbb{R}^{2N_{h}}$, defined in (2.3), is such that
(2.17) $D_{h}V=\begin{pmatrix}D_{1}\\ D_{2}\end{pmatrix}V=\begin{pmatrix}D_{1}V\\ D_{2}V\end{pmatrix}$,
where the matrices $D_{1},D_{2}\in\mathbb{R}^{N_{h}\times N_{h}}$ corresponds to the operator defined in (2.2).
For a discrete vector field $p=(p_{1},p_{2})\in\mathbb{R}^{2N_{h}}$, the discrete divergence operator ${\rm{div}}_{h}:\mathbb{R}^{2N_{h}}\to\mathbb{R}^{N_{h}}$ corresponding to (2.4) is such that
(2.18) ${\rm{div}}_{h}(p)=D_{1}p_{1}+D_{2}p_{2}=\begin{bmatrix}D_{1}&D_{2}\end{bmatrix}\begin{pmatrix}p_{1}\\ p_{2}\end{pmatrix}$.
We also denote by ${\mathop{\rm diag}}(V)\in\mathbb{R}^{N_{h}\times N_{n}}$ the diagonal matrix whose diagonal entries correspond to the components of $V$.
For $k\in\mathcal{I}^{*}_{\Delta t}$, we define the vectors $U^{n,k},U^{N_{t}}\in\mathbb{R}^{{N_{h}^{2}}}$ by
(2.19) $U^{n,k}_{i+jN_{h}}=u^{n,k}_{i,j},\qquad U^{N_{t}}_{i+jN_{h}}=G(x_{i,j})\quad\forall\,i,j\in\mathcal{I}_{h}.$
Similarly, for $k\in\mathcal{I}_{\Delta t}\setminus\{0\}$, we set $M^{n,k},M^{0}\in\mathbb{R}^{{N_{h}^{2}}}$ as
(2.20) $M^{n,k}_{i+jN_{h}}=m^{n,k}_{i,j},\qquad M^{0}_{i+jN_{h}}=m_{0}(x_{i,j})\quad\forall\,i,j\in\mathcal{I}_{h}.$
Finally, for $k\in\mathcal{I}^{*}_{\Delta t}$, we define $Q^{n,k}=(Q^{1,n,k},Q^{2,n,k})\in\mathbb{R}^{2{N_{h}^{2}}}$ as
(2.21) $Q^{1,n,k}_{i+jN_{h}}=q^{1,n,k}_{i,j},\quad Q^{2,n,k}_{i+jN_{h}}=q^{2,n,k}_{i,j}\quad\text{for all }i,j\in\mathcal{I}_{h}.$
With this notation, scheme (2.12) can be written in matrix form as
(2.22) $\begin{cases}U^{n,k}=\mathcal{A}(Q^{n,k})\,U^{n,k+1}+\Delta t\,\mathcal{W}^{k}M^{n,k}+\Delta t\,\mathcal{B}^{k},&\text{for all }k\in\mathcal{I}^{*}_{\Delta t},\\[5.69054pt] U^{n,N_{t}}=U^{N_{t}},\end{cases}$
where
(2.23) $(\mathcal{A}(Q^{n,k}))_{iN_{h}+j,p+N_{h}q}=\frac{1}{2d}\sum_{\ell=1}^{4}\beta_{p,q}(y^{\ell}(x_{i,j},q^{n,k}_{i,j})),\quad\text{for all }i,j,p,q\in\mathcal{I}_{h}$
(2.24) $(\mathcal{W}^{k})_{iN_{h}+j,p+N_{h}q}=(\operatorname{diag}\big(F^{\prime}(M^{n-1,k})\big))_{iN_{h}+j,p+N_{h}q},$
$(\mathcal{B}^{k})_{i+N_{h}j}=\frac{|q^{n,k}_{i,j}|^{2}}{2}-V(x_{i,j})$
(2.25) $\hskip 56.9055pt+F((M^{n-1,k})_{i+N_{h}j})-F^{\prime}((M^{n-1,k})_{i+N_{h}j})(M^{n-1,k})_{i+N_{h}j}\quad\text{for all }i,j\in\mathcal{I}_{h}.$
Scheme (2.15) can be written in matrix form as
(2.26) $\left\{\begin{array}[]{ll}M^{n,k+1}=\mathcal{A}(Q^{n,k})^{*}M^{n,k}+\Delta t\mathcal{Z}^{k+1}U^{n,k+1}+\Delta t\mathcal{C}^{k+1}\quad&k\in\mathcal{I}^{*}_{\Delta t},\\[6.0pt] M^{n,0}=M^{0},\end{array}\right$.
where $\mathcal{A}(Q)^{*}$ denotes the transpose of $\mathcal{A}(Q)$ given by (2.23), for every $k\in\mathcal{I}^{*}_{\Delta t}$, $\mathcal{Z}^{k+1}$ is ${N_{h}^{2}}\times{N_{h}^{2}}$ matrix such that, given $V\in\mathbb{R}^{{N_{h}^{2}}}$ and let $\mathscr{M}=M^{n-1,k+1}$
(2.27) $\mathcal{Z}^{k+1}V={\rm{div}}_{h}\big({\mathop{\rm diag}}(\mathscr{M})\,D_{h}U\big)={\mathop{\rm diag}}(\mathscr{M}){\rm{div}}_{h}(D_{h}V)+{\mathop{\rm diag}}(D_{1}\mathscr{M})D_{1}V+{\mathop{\rm diag}}(D_{2}\mathscr{M})D_{2}V.$
This shows that $\mathcal{Z}^{k+1}$ can be written in matrix form as
(2.28) $\mathcal{Z}^{k+1}={\mathop{\rm diag}}(\mathscr{M})\,L_{h}+{\mathop{\rm diag}}(D_{1}\mathscr{M})D_{1}+{\mathop{\rm diag}}(D_{2}\mathscr{M})D_{2}.$
where $L_{h}:=D_{1}D_{1}+D_{2}D_{2}$ is the discrete 2D Laplacian operator. Moreover $\mathcal{C}^{k+1}$ is the vector in $\mathbb{R}^{{N_{h}^{2}}}$
(2.29) $(\mathcal{C}^{k+1})_{i+N_{h}j}=-(\mathcal{Z}^{k+1}U^{n-1,k+1})_{i+N_{h}j}.$
The final discrete Newton iteration system reads as follows. Given $(U^{n-1,k},M^{n-1,k},Q^{n,k})_{k\in\mathcal{I}_{\Delta t}}$, find $(U^{n,k},M^{n,k})_{k\in\mathcal{I}_{\Delta t}}$ such that
(2.30) $\begin{cases}U^{n,k}=-\mathcal{A}(Q^{n,k})U^{n,k+1}+\Delta t\mathcal{W}^{k}M^{n,k}+\Delta t\mathcal{B}^{k}\quad&k\in\mathcal{I}^{*}_{\Delta t},\\[4.0pt] M^{n,k+1}=-\mathcal{A}(Q^{n,k})^{*}M^{n,k}+\Delta t\mathcal{Z}^{k+1}U^{n,k+1}+\Delta t\mathcal{C}^{k+1}\quad&k\in\mathcal{I}^{*}_{\Delta t},\\[4.0pt] U^{n,N_{t}}=U^{N_{t}},\\[4.0pt] M^{n,0}=M^{0},\end{cases}$
where $\mathcal{A}(Q^{n,k}),\;\mathcal{W}^{k},\;\mathcal{B}^{k},\;\mathcal{Z}^{k+1}$, and $\mathcal{C}^{k+1}$ are given by (2.23), (2.24), (2.4), (2.28), and (2.29) respectively.
2.5. Well-posedness
To establish the well-posedness of the system (2.30), we represent it in matrix form. For this purpose, let $\bar{U}$ and $\bar{M}$ be the vectors in $\mathbb{R}^{(N_{t}+1)N_{h}^{2}}$ defined by
(2.31) $(\bar{U}^{n})_{kN_{h}^{2}+iN_{h}+j}=(U^{n,k})_{i+N_{h}j},\quad(\bar{M}^{n})_{kN_{h}^{2}+iN_{h}+j}=(M^{n,k})_{i+N_{h}j},\quad k\in\mathcal{I}_{\Delta t},\;i,j\in\mathcal{I}_{h}.$
Next, define the block matrices $\mathbf{A}$ and $\mathbf{W}$ as
$\mathbf{A}=\begin{bmatrix}\mathcal{I}&-\mathcal{A}^{0}&\mathcal{O}&\cdots&\mathcal{O}\\[4.0pt] \mathcal{O}&\mathcal{I}&-\mathcal{A}^{1}&\ddots&\vdots\\[4.0pt] \vdots&\ddots&\ddots&\ddots&\vdots\\[4.0pt] \vdots&\ddots&\ddots&\mathcal{I}&-\mathcal{A}^{N_{t}-1}\\[4.0pt] \mathcal{O}&\cdots&\cdots&\mathcal{O}&\mathcal{I}\end{bmatrix},\qquad\mathbf{W}=\Delta t\,\begin{bmatrix}\mathcal{W}^{0}&\mathcal{O}&\cdots&\cdots&\mathcal{O}\\[4.0pt] \mathcal{O}&\mathcal{W}^{1}&\ddots&&\vdots\\[4.0pt] \vdots&\ddots&\ddots&\ddots&\vdots\\[4.0pt] \vdots&&\ddots&\mathcal{W}^{N_{t}-1}&\mathcal{O}\\[4.0pt] \mathcal{O}&\cdots&\cdots&\mathcal{O}&\mathcal{O}\end{bmatrix}$.
Here, $\mathcal{I}$ and $\mathcal{O}$ denote the $N_{h}^{2}\times N_{h}^{2}$ identity and zero matrices, respectively. If $F^{\prime}>0$ and $m>0$, then the first $N_{t}$ diagonal blocks of $\mathbf{W}$ are positive, provided that $M^{n-1}$ is positive. We also define the block diagonal matrix $\mathbf{Z}$ as
$\mathbf{Z}=\Delta t\,\begin{bmatrix}\mathcal{O}&\cdots&\cdots&\cdots&\mathcal{O}\\[4.0pt] \mathcal{O}&\mathcal{Z}^{1}&\mathcal{O}&\cdots&\vdots\\[4.0pt] \vdots&\ddots&\ddots&\ddots&\vdots\\[4.0pt] \vdots&\ddots&\ddots&\ddots&\mathcal{O}\\[4.0pt] \mathcal{O}&\cdots&\cdots&\mathcal{O}&\mathcal{Z}^{N_{t}}\end{bmatrix}$,
where, for every $k\in\mathcal{I}^{*}_{\Delta t}$, $\mathcal{Z}^{k}$ are defined in (2.28).
Finally, let
$\mathbf{B}=\Delta t\,\big[\mathcal{B}^{0},\dots,\mathcal{B}^{N_{t}-1},\frac{1}{\Delta t}U^{N_{t}}\big]^{\top},\quad\mathbf{C}=\Delta t\,\big[\frac{1}{\Delta t}M^{0},\mathcal{C}^{0},\dots,\mathcal{C}^{N_{t}-1}\big]^{\top},$
with $\mathcal{B}^{k}$ and $\mathcal{C}^{k}$ given in (2.4) and (2.29), respectively.
Then, the discrete Newton system (2.30) can be written as a Hamiltonian system:
(2.32) $\begin{bmatrix}\mathbf{A}&-\mathbf{W}\\[5.69054pt] -\mathbf{Z}&-\mathbf{A}^{\top}\end{bmatrix}\begin{bmatrix}\bar{U}^{n}\\[2.84526pt] \bar{M}^{n}\end{bmatrix}=\begin{bmatrix}\mathbf{B}\\[2.84526pt] \mathbf{C}\end{bmatrix}$.
Proposition 2.1.
Assume that $F^{\prime}>0$ and $\bar{M}^{n-1}>0$ for any $n\in\mathbb{N}$. Then, there exists a unique solution $(\bar{U}^{n},\bar{M}^{n})$ to the system (2.32).
Proof.
Suppose that ${\textbf{B}}=0$ and ${\textbf{C}}=0$, then (2.32) reads as
(2.33) $\left\{\begin{array}[]{l}\textbf{A}\,{\bar{U}^{n}}-\textbf{W}{\bar{M}^{n}}=0,\\ -\textbf{Z}\,{\bar{U}^{n}}-\textbf{A}^{\top}{\bar{M}^{n}}=0\end{array}\right$.
Multiplying the first equation by $(\bar{M}^{n})^{\top}$ and the second one by $(\bar{U}^{n})^{\top}$ one gets
$\left\{\begin{array}[]{l}(\bar{M}^{n})^{\top}\textbf{A}\,{\bar{U}^{n}}-(\bar{M}^{n})^{\top}\textbf{W}{\bar{M}^{n}}=0\\ -(\bar{U}^{n})^{\top}\textbf{Z}\,{\bar{U}^{n}}-(\bar{U}^{n})^{\top}\textbf{A}^{\top}{\bar{M}^{n}}=0.\end{array}\right$.
Adding both equations, we obtain
(2.34) $(\bar{M}^{n})^{\top}\textbf{W}{\bar{M}^{n}}+(\bar{U}^{n})^{\top}\textbf{Z}\,{\bar{U}^{n}}=0.$
Since $F^{\prime}>0$, the block $\mathcal{W}^{k}$ is positive definite for all $k\in\mathcal{I}_{\Delta t}$. Moreover, the block $\mathcal{Z}^{k}$ is positive definite for all $k\in\mathcal{I}_{\Delta t}$, since $\bar{M}^{n-1}>0$ and it is the sum of a positive definite matrix and a skew symmetric matrix and then, by [J, Remark 1], it is positive definite.
Hence, it follows from (2.34) that
[…]
and
[…]
Finally, replacing (2.35) and (2.36) in (2.33), we get $\bar{U}^{n}=0$ and $\bar{M}^{n}=0$, and then the existence of a unique solution to (2.32). ∎
Remark 2.1.
Le us mention that the assumption $F^{\prime}>0$ is a standard monotonicity assumption in MFG with separate Hamiltonian, which guarantees the uniqueness of solution to (1.1), see [newton, Remark 2.1]. However, in some practical applications, the semi-Lagrangian scheme can still be applied successfully even when these assumptions are not satisfied, as illustrated in the numerical tests in Section 4.
Remark 2.2.
Note that the blocks of W would be dense matrices if $F(m(t,x))$ is replaced by a nonlocal operator $f[m(t,\cdot)]$. In this case, we use the notation $\frac{\delta f}{\delta m}$ for the flat derivative of $f$ (see e.g [mfgnash]). The condition $F^{\prime}>0$ can then be replaced by assuming that $\frac{\delta f}{\delta m}$ is Lipschitz continuous and that $f$ satisfies
$\int_{\mathbb{T}^{d}}(f[m](x)-f[m^{\prime}](x))\,{\rm d}(m-m^{\prime})(x)$ $\geq 0\quad\text{for all }m,m^{\prime}.$
A typical example is a nonlocal coupling with smoothing effect
$f(x,m)=\int_{\mathbb{T}^{d}}\Phi(z,(\rho\ast m)(z))\rho(x-z){\rm d}z,$
where $\ast$ denotes the usual convolution on $\mathbb{T}^{d}$, $\Phi$ is a smooth and nondecreasing with respect to $m$, and $\rho$ is a smooth, even function with compact support. In this case, writing $\Phi=\Phi(x,\theta)$, the flat derivative of $f$ reads
$\frac{\delta f}{\delta m}(x,m,y)=\int_{\mathbb{T}^{d}}\frac{\partial\Phi}{\partial\theta}(z,(\rho\ast m)(z))\rho(x-z)\rho(z-y){\rm d}z.$
3. A finite differences scheme
In this section, we discretize the iterative system (1.5) using an implicit finite difference scheme in two dimensions and prove its well-posedness. The equation for $v$ is discretized using upwind differences for the drift term and central differences for the second-order spatial derivatives. The equation for $m$ is then approximated by taking the adjoint of the resulting scheme.
In order to discretize the system, we define the discrete time derivative of $v\in B(\mathcal{G}_{\Delta t}\times\mathcal{G}_{h})$ as
(3.1) $D_{t}v^{k}_{i,j}=\frac{v^{k+1}_{i,j}-v^{k}_{i,j}}{\Delta t}\quad\text{for all }k\in\mathcal{I}_{\Delta t}^{*},\;i,j\in\mathcal{I}_{h},$
Given $q^{n}\in B(\mathcal{G}_{\Delta t}\times\mathcal{G}_{h})^{2}$ and $u^{n-1},m^{n-1},m^{n-1}\in B(\mathcal{G}_{\Delta t}\times\mathcal{G}_{h})$, we discretize the first equation in (1.5) using an implicit scheme finite difference scheme, which reads for all $k\in\mathcal{I}_{\Delta t}^{*}$ and $i,j\in\mathcal{I}_{h}$ as
(3.2) $\begin{cases}\displaystyle-D_{t}u^{n,k}_{i,j}-\nu(L_{h}u^{n,k})_{i,j}+q^{n,k}_{i,j}(D_{h}u^{n,k})_{i,j}=&\frac{1}{2}|(q^{n,k})_{i,j}|^{2}+V(x_{i,j})+\mathcal{F}_{i,j}(m^{n,k+1},m^{n-1,k+1}),\\[6.0pt] u^{n,T}_{i,j}=G(x_{i,j}),\end{cases}$
where $L_{h}$ is thediscrete Laplace operator, already defined in Section 2.4, and for every $\zeta,\tilde{\zeta}\in B(\mathcal{G}_{h})$
$\mathcal{F}_{i,j}(\zeta,\tilde{\zeta})=F^{\prime}(\tilde{\zeta}_{i,j})(\zeta_{i,j}-\tilde{\zeta}_{i,j})+F(\tilde{\zeta}_{i,j}).$
Then, we consider the following approximation of the forward equation in (1.5)
(3.3) $\begin{cases}D_{t}m^{n,k}_{i,j}-\nu(L_{h}m^{n,k+1})_{i,j}-({\rm{div}}_{h}(q^{n,k}m^{n,k+1}))_{i,j}={\rm{div}}_{h}(m^{n-1,k+1}(D_{h}u^{n,k}-D_{h}u^{n-1,k}))_{i,j},\\[6.0pt] m^{n,0}_{i,j}=m_{0}(x_{i,j}).\end{cases}$
Remark 3.1.
Notice that, qiven $q\in B(\mathcal{G}_{\Delta t}\times\mathcal{G}_{h})^{2}$, the operator $\mathcal{B}(\mathcal{G}_{h})\ni m\to(-\nu(L_{h}m)-({\rm{div}}_{h}(qm))\in\mathcal{B}(\mathcal{G}_{h})$ is the adjoint of the operator $\mathcal{B}(\mathcal{G}_{h})\ni u\to(-\nu(L_{h}u)+q\,D_{h}u)\in\mathcal{B}(\mathcal{G}_{h})$.
Combining (3.2) and (3.3), and using the vector notations (2.19)-(2.21), we get the following fully discrete upwind scheme for the Newton iterations system (1.5)
(3.4) $\begin{cases}\mathcal{D}^{k}U^{n,k}=U^{n,k+1}+\Delta t\mathcal{W}^{k+1}M^{n,k+1}+\Delta t\tilde{\mathcal{B}}^{k+1}\quad&\text{for }k\in\mathcal{I}^{*}_{\Delta t},\\[4.0pt] (\mathcal{D}^{k})^{\top}M^{n,k+1}=M^{n,k}+\Delta t\mathcal{Z}^{k}U^{n,k}+\Delta t\tilde{\mathcal{C}}^{k+1}\quad&\text{for }k\in\mathcal{I}^{*}_{\Delta t},\\[4.0pt] U^{n,N_{t}}=U^{T},\\[4.0pt] M^{n,0}=M^{0},\end{cases}$
where $\mathcal{D}^{k}$ is the ${N_{h}^{2}}\times{N_{h}^{2}}$ matrix
$\mathcal{D}^{k}=\mathcal{I}-\Delta t(\nu L_{h}+Q_{1}D_{1}+Q_{2}D_{2}),$
and $\tilde{\mathcal{C}}^{k+1},\tilde{\mathcal{B}}^{k+1}$ are vectors in $\mathbb{R}^{N_{h}^{2}}$ defined as
$\begin{split}(\tilde{\mathcal{C}}^{k+1})_{i+N_{h}j}&=-(\mathcal{Z}^{k+1}U^{n-1,k})_{i+N_{h}j},\\ (\tilde{\mathcal{B}}^{k+1})_{i+N_{h}j}&=\frac{|(q^{n,k})_{i,j}|^{2}}{2}-V(x_{i,j})\\ &\;+F((M^{n-1,k+1})_{i+N_{h}j})-F^{\prime}((M^{n-1,k+1})_{i+N_{h}j})(M^{n-1,k+1})_{i+N_{h}j}\quad\text{for }i,j\in\mathcal{I}_{h}.\end{split}$
Finally, define $\tilde{U}$ and $\tilde{M}\in\mathbb{R}^{(N_{t})N_{h}^{2}}$ by
$\begin{split}(\tilde{U}^{n})_{kN_{h}^{2}+iN_{h}+j}&=(U^{n,k})_{i+N_{h}j}\quad\text{for }k\in\mathcal{I}^{*}_{\Delta t},\,i,j\in\mathcal{I}_{h},\\ (\tilde{M}^{n})_{kN_{h}^{2}+iN_{h}+j}&=(M^{n,k})_{i+N_{h}j}\quad\text{for }k\in\mathcal{I}_{\Delta t}\setminus\{0\},\,i,j\in\mathcal{I}_{h}.\end{split}$
In order to write system (3.4) in a matrix way, as in (2.32), we define first the matrices
$\tilde{\textbf{A}}=\left[\begin{matrix}\mathcal{D}^{0}&-\mathcal{I}&\mathcal{O}&\cdots&\mathcal{O}\\ \mathcal{O}&\mathcal{D}^{1}&\ddots&\ddots&\vdots\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ \vdots&\ddots&\ddots&\ddots&-\mathcal{I}\\ \mathcal{O}&\cdots&\cdots&\mathcal{O}&\mathcal{D}^{N_{t}-1}\end{matrix}\right],\quad\tilde{\textbf{W}}=\Delta t\left[\begin{matrix}\mathcal{W}^{1}&\mathcal{O}&\cdots&\cdots&\mathcal{O}\\ \vdots&\mathcal{W}^{2}&\ddots&\ddots&\mathcal{O}\\ \mathcal{O}&\ddots&\ddots&\ddots&\vdots\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ \mathcal{O}&\cdots&\cdots&\cdots&\mathcal{W}^{N_{t}}\\ \end{matrix}\right],$
and
$\tilde{\textbf{Z}}=\Delta t\left[\begin{matrix}\mathcal{Z}^{0}&\mathcal{O}&\cdots&\cdots&\mathcal{O}\\ \mathcal{O}&\mathcal{Z}^{1}&\ddots&\ddots&\mathcal{O}\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ \mathcal{O}&\mathcal{O}&\cdots&\cdots&\mathcal{Z}^{N_{t}-1}\\ \end{matrix}\right].$
Hence, (3.4) is equivalent to the Hamiltonian system
(3.5) $\left[\begin{matrix}\tilde{\textbf{A}}&-{\tilde{\textbf{W}}}\\[6.0pt] -{\tilde{\textbf{Z}}}&-\tilde{\textbf{A}}^{*}\end{matrix}\right]\left[\begin{matrix}\tilde{U}^{n}\\[6.0pt] \tilde{M}^{n}\end{matrix}\right]=\left[\begin{matrix}\tilde{\textbf{B}}\\[6.0pt] \tilde{\textbf{C}}\end{matrix}\right$],
where $\tilde{\textbf{B}}=\Delta t\Big[\mathcal{B}^{1},\dots,\mathcal{B}^{N_{t}}+\frac{1}{\Delta t}U^{N_{t}}\Big]^{\top}$ and $\tilde{\textbf{C}}=\Delta t\Big[\frac{1}{\Delta t}M^{0}+\mathcal{C}^{1},\dots,\mathcal{C}^{N_{t}}\Big]^{\top}$. Then, arguing as in the proof of Proposition 2.1, one can show the following well-posedness result.
Proposition 3.1.
Suppose that $\tilde{M}^{n-1}>0$ for any $n\in\mathbb{N}$. Then, there exists a unique solution $(\tilde{U}^{n},\tilde{M}^{n})$ to the system (3.5).
4. Numerical tests
In this section, we assess the performance of the proposed numerical schemes through a series of tests in both one and two spatial dimensions, considering cases with separable and non-separable Hamiltonians. In all tests, the Newton iterations are terminated when the following error metrics fall below a prescribed tolerance $\tau$, set to $10^{-4}$ in all the experiments. Specifically, we monitor the relative variation between two consecutive Newton iterates for both the value function and the distribution, according to:
(4.1) $E(m^{n}):=\|m^{n+1}-m^{n}\|_{\infty}\quad\text{and}\quad E(u^{n}):=\|u^{n+1}-u^{n}\|_{\infty}.$
The systems defined by equations (2.32) and (3.5) are solved using Gauss–Seidel iterations. The iterations are terminated once the uniform norms of the differences between consecutive solutions fall below a prescribed threshold $\delta$, which is set to $10^{-4}$. This ensures comparable accuracy for both discretization procedures. We compare three approaches for solving system (1.1): the semi-Lagrangian scheme (2.30) (Newton–SL), the finite difference scheme (3.4) (Newton–FD), and the finite difference schemes from [finitedifference1] solved via Newton iterations (FD–Newton). Algorithm 1 implements the Newton method for both Newton–SL and Newton–FD. We note that the semi-Lagrangian scheme used in this work is explicit and does not require a parabolic CFL condition, whereas the finite difference schemes are implicit and thus impose no restriction on the time step. For accuracy purposes, the semi-Lagrangian scheme is run with a time step of the form $\Delta t=O(h^{3/2})$ (see [Ferretti1, Ferretti2] for a detailed analysis). We now present a series of four numerical tests designed to illustrate the behavior and performance of the proposed Newton-based schemes. The first test considers a one-dimensional MFG with a known reference solution, allowing us to validate the accuracy of the schemes and compare the performance of Newton–SL, Newton–FD, and FD–Newton. The second test focuses on one-dimensional MFG with smooth initial and terminal conditions, considered for two different diffusion coefficients ($\nu=0.4$ and $\nu=0.02$), highlighting the stability and robustness of the schemes. In the small-diffusion case ($\nu=0.02$), we observe a break-down of the standard Newton iterations for the finite-difference schemes (Newton–FD and FD–Newton) due to a poor initialization, whereas Newton–SL remains stable. To overcome this difficulty, we employ a globalized version of Newton’s method, which is activated whenever the local iteration fails; this global strategy is described in detail in the following subsection. Given the robust and accurate performance of the Newton–SL scheme observed in the tests 1–2, the subsequent experiments involving non-separable Hamiltonian are conducted using only this discretization approach. The third test examine one-dimensional MFG system with non-separable Hamiltonian, and finally the fourth test examine a two-dimensional MFG systems with a separable Hamiltonian. Both tests demonstrate the flexibility and robustness of the Newton–SL scheme in handling more general and nonlinear Hamiltonian structures. Together, these tests provide a comprehensive evaluation of the schemes, highlighting their accuracy, efficiency, and stability across different problem types, dimensions, and diffusion regimes.
Algorithm 1 Local Newton iterations for mean field games
1:Input: Initial guesses $u^{0}$, $m^{0}$, and tolerance $\tau$
[…]
4:repeat
5:  Compute $(m^{n+1},u^{n+1})$ as the solution of (2.30) (Newton–SL) or (3.4) (Newton–FD)
6:  $E(m^{n})\leftarrow\|m^{n+1}-m^{n}\|_{\infty}$
7:  $E(u^{n})\leftarrow\|u^{n+1}-u^{n}\|_{\infty}$
8:  Update $q^{n}\leftarrow D_{h}u^{n}$
9:  $n\leftarrow n+1$
10:until $E(m^{n})>\tau$ or $E(u^{n})>\tau$
11:Output: Approximated solution $(u^{n+1},m^{n+1})$ to the MFG system (1.1)
Remark 4.1.
For the semi-Lagrangian scheme, we approximate the drift term $q^{n}$ using the following discrete formulation:
$q^{n,k}_{i,j}=\left(D\rho_{\varepsilon}*u^{n-1,k}\right)_{i,j}\quad\text{for all }k\in\mathcal{I}_{\Delta t},\,i,j\in\mathcal{I}_{h},$
where * denotes the discrete circular convolution between two grid functions $A,B\in\mathbb{R}^{N_{h}\times N_{h}}$ (see e.g [oppenheim2010discrete, Chapter 8]), and, for $\varepsilon>0,$ the kernel $\rho_{\varepsilon}$ is a Gaussian mollifier defined by
$\rho_{\varepsilon}(x)=\frac{1}{\sqrt{2\pi}\varepsilon}e^{-\frac{|x|^{2}}{2\varepsilon^{2}}},\quad\text{for all }x\in\mathbb{R}^{d}.$
This approach, which is common in the framework of SL schemes (e.g see [CSorder2, carlini2023lagrangegalerkin]), enhances the accuracy of the scheme and allows us to avoid the constraints of a parabolic CFL condition, which would otherwise limit the time step. For the finite difference scheme, we use instead the direct formulation:
$q^{n,k}_{i,j}=(D_{h}u^{n-1,k})_{i,j}\quad\text{for all }k\in\mathcal{I}_{\Delta t},\,i,j\in\mathcal{I}_{h},$
where $D_{h}$ is given by (2.3).
4.1. Global Newton algorithm
In this section, we describe the global Newton method, which extends the local convergence properties of the standard Newton method to a global setting. The key idea behind this approach is to modify the Newton step when the current iterate is far from the solution, ensuring that each iteration makes sufficient progress. The global convergence of the Newton method in infinite dimensions is proved in [global-newton]. We adapt this framework to our MFG problem to enhance robustness and guarantee convergence even when starting from rough or distant initial guesses. First, we define the search direction $d^{n}:=(v^{n},\rho^{n})$ as the solution to the linear system
(4.2) $J{\mathcal{F}}(u^{n},m^{n})d^{n}=-{\mathcal{F}}(u^{n},m^{n}).$
Then, the globalized Newton method generates sequences $\{(u^{n},m^{n})\}$, $\{d^{n}\}$, and $\{\alpha^{n}\}$, related by the iteration:
$(u^{n+1},m^{n+1})=(u^{n},m^{n})+\alpha^{n}d^{n},\quad n\in\mathbb{N},$
where the step length $\alpha^{n}>0$ is determined by a line-search procedure using a suitably defined merit function. Following [global-newton], we use the squared $L^{2}$-norm of $\mathcal{F}$ as the merit function:
(4.3) $\begin{array}[]{rcl}\displaystyle\Theta(u,m):=\frac{1}{2}\|\mathcal{F}(u,m)\|_{2}^{2}&=&\int_{[0,T]\times{\mathbb{T}}^{d}}|-\partial_{t}u-\nu\Delta u+H(x,Du)-F(m(t,x))|^{2}{\rm d}x{\rm d}t\\[10.0pt] &+&\int_{[0,T]\times{\mathbb{T}}^{d}}|\partial_{t}m-\nu\Delta m-{\rm{div}}(mD_{p}H(x,Du)|^{2}{\rm d}x{\rm d}t\\[10.0pt] &+&\int_{\mathbb{T}^{d}}|u(T,x)-G(x)|^{2}{\rm d}x+\int_{\mathbb{T}^{d}}|m(0,x)-m_{0}(x)|^{2}{\rm d}x.\end{array}$
A reasonable solution for the system should correspond to a root of the merit function, i.e., $\Theta(u,m)=0$, which implies that $\mathcal{F}(u,m)=0$ is satisfied, meaning that the system of equations is fulfilled. The step length $\alpha^{n}$ is determined by enforcing the Armijo condition, which guarantees a sufficient decrease of the merit function at each iteration. Specifically, this condition requires
$\Theta(u^{n+1},m^{n+1})\leq\Theta(u^{n},m^{n})+c\alpha^{n}\,\Theta^{\prime}(u^{n},m^{n})(d^{n}),$
where $0<c<1/2$ is a constant controlling the sufficient decrease, and $\Theta^{\prime}$ denotes the Fréchet derivative of $\Theta$. Moreover, when the direction $d^{n}$ satisfies the Newton equation (4.2), it holds that
(4.4) $\Theta^{\prime}(u^{n},m^{n})d^{n}=-2\,\Theta(u^{n},m^{n})=-\|\mathcal{F}(u^{n},m^{n})\|^{2}.$
The global Newton procedure for solving (1.1) is summarized in Algorithm 2. Notice that the Newton step (4.2) leads to the linear parabolic system for $d^{n}=(v^{n},\rho^{n})$, as given in (1.5). At the discrete level, the search direction $d^{n}$ can be computed using either the Newton-SL or the Newton-FD scheme.
Algorithm 2 Global Newton iterations for mean field games
1:Input: Initial guesses $u^{0}$, $m^{0}$, parameters $\beta\in(0,1)$, $c\in(0,\frac{1}{2})$, and tolerance $\tau>0$
[…]
4:repeat
5:  Compute $d^{n}=(m^{n+1/2},u^{n+1/2})$ s the solution of (2.30) (Newton–SL) or (3.4) (Newton–FD)
6:  Find the smallest integer $i_{n}\geq 0$ such that the step length $\alpha_{n}=\beta^{i_{n}}$ satisfies the Armijo condition:
$\Theta((m^{n},u^{n})+\alpha_{n}d^{n})\leq\Theta(m^{n},u^{n})+c\alpha^{n}\Theta^{\prime}(m^{n},u^{n})d^{n}$
7:  Update the iteration $(m^{n+1},u^{n+1})\leftarrow(m^{n},u^{n})+\alpha_{n}d^{n}$
8:  $E(m^{n})\leftarrow\|m^{n+1}-m^{n}\|_{\infty}$
9:  $E(u^{n})\leftarrow\|u^{n+1}-u^{n}\|_{\infty}$
10:  Update $q^{n}\leftarrow D_{h}u^{n}$
11:  $n\leftarrow n+1$
12:until $E(m^{n})>\tau$ or $E(u^{n})>\tau$
13:Output: Approximated solution $(u^{n+1},m^{n+1})$ to the MFG system (1.1)
Remark 4.2.
The globalized Newton method typically requires more iterations than the local Newton method. Therefore, in our numerical experiments, we first apply the local Newton iterations, and only if they fail to converge, we switch to the globalized version.
4.2. Test 1: Accuracy Comparison for an MFG Problem with Reference Solution
In this test (see [popov]), we consider the MFG system (1.1) on the time-space domain $0,0.05]\times]0,1[$ with periodic boundary conditions, $\nu=0.05$, and $H(x,p)=\frac{|p|^{2}}{2}$. The coupling term is giving by
$F(m(x))=4\min(4,m)-3m_{0}(x),\quad\text{for all }x\in]0,1[,$
while the initial condition is given by
$m_{0}(x)=\begin{cases}4\sin^{2}(2\pi(x-{1}/4))\quad&\text{if }x\in[{1}/4,{3}/4],\\[6.0pt] 0&\text{otherwise},\end{cases}$
and the terminal condition $G$ is set to to zero. To evaluate the accuracy errors, we compare the approximate solutions with a reference solution computed using a high-order value iteration scheme [Calzola_Carlini_Silva_lg_second_order] on a fine mesh with spatial step $h=6.67\times 10^{-4}$ and time step $\Delta t=h^{3/2}$. The accuracy of each scheme is assessed by computing the error in the discrete uniform norm. In our simulations, we set $\Delta t=h^{3/2}/2$ for the Newton–SL method, and $\Delta t=h/4$ for both Newton–FD and FD–Newton. It is worth noting that the semi-Lagrangian scheme is unconditionally stable, and the chosen time step is primarily dictated by accuracy considerations; in particular, we adopt the same time step used in [Calzola_Carlini_Silva_lg_second_order], from which the reference solution is taken. On the other hand, the finite-difference scheme is implicit, allowing for relatively large time steps without any CFL restriction, so that the time step selection in this case is also driven by accuracy rather than stability constraints.
Table 1 reports the numerical results obtained for different grid resolutions $h$, including the errors, CPU times, and number of Newton iterations for the three methods under consideration. The discrete uniform errors are defined as
$E_{\infty}(u)=\|u^{n}-u_{\mathrm{ref}}\|_{\infty},\qquad E_{\infty}(m)=\|m^{n}-m_{\mathrm{ref}}\|_{\infty},$
where $(u_{\mathrm{ref}},m_{\mathrm{ref}})$ denotes the reference solution, and $(u^{n},m^{n})$ denotes the numerical solution obtained by the Newton method at the iteration when the stopping criterion is satisfied.
We first observe that the Newton–SL approach requires fewer iterations and consistently less computational time compared to the two finite-difference-based schemes, while achieving an accuracy comparable to theirs. Between FD–Newton and Newton–FD, the results are very close in terms of both accuracy and CPU time, with Newton–FD showing a slight advantage in efficiency.
Newton-SL with $\Delta t=h^{3/2}/2$
$h$ $E_{\infty}(u)$ $E_{\infty}(m)$ Time Iterations
2.50 $\cdot 10^{-2}$ 5.51 $\cdot 10^{-2}$ 1.64 $\cdot 10^{-1}$ 0.61s 6
1.25 $\cdot 10^{-2}$ 2.40 $\cdot 10^{-2}$ 1.16 $\cdot 10^{-1}$ 2.77s 7
6.25 $\cdot 10^{-3}$ 1.83 $\cdot 10^{-2}$ 6.61 $\cdot 10^{-2}$ 13.92s 7
3.125 $\cdot 10^{-3}$ 4.50 $\cdot 10^{-3}$ 1.41 $\cdot 10^{-2}$ 80.60s 7
FD-Newton with $\Delta t=h/4$
$h$ $E_{\infty}(u)$ $E_{\infty}(m)$ Time Iterations
2.50 $\cdot 10^{-2}$ 1.23 $\cdot 10^{-1}$ 3.11 $\cdot 10^{-2}$ 2.23s 7
1.25 $\cdot 10^{-2}$ 6.21 $\cdot 10^{-2}$ 1.63 $\cdot 10^{-2}$ 18.32s 8
6.25 $\cdot 10^{-3}$ 3.14 $\cdot 10^{-2}$ 8.75 $\cdot 10^{-3}$ 92.91s 8
3.125 $\cdot 10^{-3}$ 1.77 $\cdot 10^{-2}$ 9.54 $\cdot 10^{-3}$ 597.21s 8
Newton-FD with $\Delta t=h/4$
$h$ $E_{\infty}(u)$ $E_{\infty}(m)$ Time Iterations
2.50 $\cdot 10^{-2}$ 1.532 $\cdot 10^{-1}$ 3.42 $\cdot 10^{-2}$ 1.48s 7
1.25 $\cdot 10^{-2}$ 6.71 $\cdot 10^{-2}$ 1.83 $\cdot 10^{-2}$ 12.27s 7
6.25 $\cdot 10^{-3}$ 3.37 $\cdot 10^{-2}$ 9.51 $\cdot 10^{-3}$ 68.10s 7
3.125 $\cdot 10^{-3}$ 1.91 $\cdot 10^{-2}$ 7.38 $\cdot 10^{-3}$ 436.01s 7
Table 1. Test 1. Comparison of Newton–SL, FD–Newton, and Newton–FD schemes: errors, CPU times, and number of iterations for the solution $(u,m)$.
4.3. Test 2: a MFG problem varying the diffusion
In this test, we consider two MFG systems, also numerically studied in [litest], defined on the time–space domain $0,0.01]\times(0,1)$ with periodic boundary conditions and diffusion coefficients $\nu=0.4$ and $\nu=0.02$, respectively. The Hamiltonian is given by $H(x,p)=\tfrac{|p|^{2}}{2}-V(x)$, and the corresponding data, illustrated in Figure 1, are
$\begin{array}[]{rcl}m_{0}(x)&=&1+\frac{1}{2}\cos(2\pi x),\\ G(x)&=&\sin(4\pi x)+0.1\cos(10\pi x),\\ V(x)&=&200\cos(2\pi x)-10\cos(4\pi x),\\ F(m)&=&m^{2}.\end{array}$
The purpose of this test is to assess the performance of the proposed methods in both a highly diffusive regime and a nearly deterministic setting, and to investigate how the diffusion parameter influences the convergence behavior and stability of the Newton iterations. The spatial and temporal discretization parameters are set to $h=6.25\cdot 10^{-3}$ and $\Delta t=h/4$ for the two finite-difference schemes, while $\Delta t=h^{3/2}/2$ is used for the Newton–SL method.
(a) Initial distribution $m_{0}$
(b) Terminal cost $G$
(c) Potential $V$
Figure 1. Test 2. The initial data.
4.3.1. Test 2.1 with $\nu=0.4$
We first consider the case $\nu=0.4$, and solve the MFG system using the three schemes: Newton–SL, Newton–FD, and FD–Newton.
(a) $E(m^{n})$
(b) $E(u^{n})$
Figure 2. Test 2.1. Convergence history of the Newton iterations for Newton–SL, Newton–FD, and FD–Newton.
Figure 3. Log–log plot of $\log(m^{n})$ versus $\log(m^{n+1})$ for the three Newton-type schemes (Newton–SL, Newton–FD and FD–Newton).
(a) Newton-SL
(b) Newton-FD
(c) FD-Newton
Figure 4. Test 2.1. Space–time evolution of the approximated distributions $m$ computed by Newton–SL, Newton–FD, and FD–Newton.
(a) Newton-SL
(b) Newton-FD
(c) FD-Newton
Figure 5. Test 2.1. Space–time evolution of approximated value function $u$ computed by Newton-SL, Newton-FD, and FD-Newton methods.
Figure 2 shows, in logarithmic scale, the evolution of the Newton iteration errors $E(m^{n})$ and $E(u^{n})$ for the three schemes: Newton–SL, Newton–FD, and FD–Newton. All the three schemes exhibit comparable convergence behavior, whereas FD–Newton shows a slightly faster decay of the $\ell^{\infty}$–norm of successive iterates.
In Figure 3 we compare the convergence rate of the three Newton algorithms by plotting $\log(m^{n+1})$ against $\log(m^{n})$. The estimated convergence rate is computed as the slope of the least-squares linear fit of the data points $(\log(m^{n}),\log(m^{n+1}))$, which provides a numerical estimate of the exponent $p$ in the relation $m^{n+1}\approx C(m^{n})^{p}$. This results in $p=1.408$ for Newton–SL, $p=1.393$ for Newton–FD and $p=1.663$ for FD–Newton. For a quadratically exact convergent method we expect $p=2$, hence an ideal slope of $2$ in this log–log plot.
Figure 4 displays the space–time evolution of the approximated distributions computed by the three schemes. It can be observed that, at the final time, the density concentrates around the points where the terminal cost $G$, shown in Figure 1(b), attains its minima. Finally, Figure 5 presents the corresponding space–time plots of the approximated value functions obtained with the same schemes. The solutions appear smooth, as expected for the relatively large diffusion coefficient $\nu=0.4$. Overall, the three methods produce visually consistent results, in good agreement with the reference test reported in [litest], confirming the reliability and coherence of the proposed discretizations.
4.3.2. Test 2.2
We now consider a smaller diffusion coefficient, specifically $\nu=0.02$. As shown in Figure 6(a), Newton–SL iterations converge after seven steps, reaching the prescribed tolerance. In contrast, both Newton–FD and FD–Newton iterations experience breakdown after only a few iterations, as the computed distribution becomes negative. This behavior highlights the increased robustness of the Newton–SL scheme in regimes characterized by low diffusion. Figures 6(b)6(c) display the space–time evolution of the approximated distribution and the corresponding value function obtained with the three methods. Since both FD–Newton and Newton–FD fail when applying local iterations, we employ the global Newton method described in Section 4.1, using the parameters $\beta=1/2$ and $c=1/3$ for the line-search procedure. These values ensures sufficient decrease of the merit function and robust globalization of the iterations. As expected, both schemes converge under this approach. Figure 7 shows the Newton residual error together with the time evolution of the approximated distribution and value function obtained with the Newton–FD scheme corresponding to the global Newton iterations. For brevity, we omit the analogous results for FD–Newton, whose performance is comparable to that of Newton–FD. These results confirm the effectiveness of the globalized Newton strategy in stabilizing the discretizations and achieving convergence even in challenging small-diffusion MFG regimes.
Remark 4.3.
In [Achlau], the authors solve a finite difference discretization of the MFG system by employing Newton’s method combined with a continuation strategy with respect to the diffusion parameter $\nu$ (see also [achdou1, achdou2]). This approach is particularly effective for handling problems with small diffusion values. Specifically, the problem is first solved for a large value of $\nu$, and the corresponding solution is then used as the initial guess to solve, still by Newton’s method, the discrete MFG system with a smaller diffusion coefficient. The procedure is repeated until the target (small) viscosity is reached. We point out that such a continuation approach has a similar effect to the global Newton iterations, as both techniques aim to improve the robustness of convergence when the local Newton method fails due to poor initialization.
(a) Convergence history of the Newton iterations for Newton–SL.
(b) Evolution of $m$
(c) Evolution of $u$
Figure 6. Convergence history of the Newton–SL iterations and space–time evolution of the approximated distribution $m$ and value function $u$ for $\nu=0.02$
(a) Convergence history for the global Newton iterations for Newton–FD.
(b) Evaluation of $m$
(c) Evaluation of $u$
Figure 7. Test $2.$ Convergence history of the global Newton-FD iterations and space–time evolution of the approximated distribution $m$ and value function $u$ for $\nu=0.02$.
4.4. Test 3: a MFG problem with non separable Hamiltonian
We consider a one-dimensional MFG system featuring a non-separable Hamiltonian,
$\begin{cases}&-\partial_{t}u-0.05\Delta{u}+\frac{1}{2(1+4m)^{\gamma}}\lvert Du\rvert^{2}=\zeta m\quad\text{in }[0,1]\times[0,1],\\ &\partial_{t}m-0.05\Delta m-\text{div}\left(\frac{mDu}{(1+4m)^{\gamma}}\right)=0\quad\text{in }[0,1]\times[0,1],\\ &G(x)=10\min\{(x-0.3)^{2},\,(x-0.7)^{2}\}\quad\text{in }[0,1],\\ &m_{0}(x)=4\;\text{for }x\in[0.375,0.625],\;m_{0}(x)=0\;\text{otherwise}\end{cases}$
In this test, we compute the solution of the system using the Newton–SL scheme, with parameters $\gamma=1.5$, $\zeta=1$, spatial step $h=5\cdot 10^{-3}$, and time step $\Delta t=h^{3/2}/2$.
In Figure 8, we display the time evolution of the approximated density and value function at the time instants $t_{k}\in\{0,0.2,0.4,0.6,0.8,1\}$. We observe that the population distribution splits into two groups: one moving toward the left target and the other toward the right target. However, due to the presence of congestion and crowd-aversion effects, neither group fully concentrates on its respective target at the final time.
Figure 8. Test 3. Time evolution of the approximated density (left) and approximated value function (right)
For this non-separable Hamiltonian test, the simulation obtained with the Newton–SL scheme is stable and accurate, and remains fully consistent with the reference outcomes reported in [litest].
4.5. Test 4: a MFG with separable Hamiltonian in dimension two
We now consider a two-dimensional MFG system, previously investigated numerically in [litest]. The computational domain is $0,1]\times(0,1)^{2}$, with diffusion coefficient $\nu=0.4$, and the following data:
$m_{0}(x_{1},x_{2})$ $=1+12\cos(2\pi x_{1})+12\cos(2\pi x_{2}),$
$G(x_{1},x_{2})$ $=\cos(2\pi x_{1})+\cos(2\pi x_{2}),$
$H(x_{1},x_{2},p)$ $=|p|^{2}+\sin(2\pi x_{1})+\sin(2\pi x_{2})+\cos(4\pi x_{1}),$
$F(m)$ $=m^{2}.$
We set $h=10^{-2}$ and $\Delta t=h^{3/2}/2$. The evolution of the approximated distribution $m$, computed with the Newton–SL scheme, is displayed in Figure 9 at times $t_{k}\in\{0,0.5,0.75,1\}$. Figures 9(b)9(c) illustrate the stationary profile attained by the density at intermediate times, which can be interpreted as a manifestation of the turnpike effect. Turnpike phenomena for mean field games with local couplings have been discussed in [turnpike]. The results are consistent with those reported in [litest], confirming the accuracy of the proposed scheme in reproducing two-dimensional MFG dynamics.
(a) $t_{k}=0$
(b) $t_{k}=N_{t}/2$
(c) $t_{k}=3N_{t}/4$
(d) $t_{k}=N_{t}$
Figure 9. Test 4. Time evolution of the approximated distribution $m$ at with Newton-SL at different times.
5. Conclusion
In this work, we have proposed a new technique for numerically solving a second-order mean field games system with local coupling by applying Newton iterations in infinite dimensions. Our study is based on a series of numerical experiments comparing two possible strategies: performing Newton iterations at the continuous level and then discretizing, or discretizing first and applying Newton iterations to the resulting finite-dimensional non linear system. A first remark is that the former strategy leads to a relatively simple scheme, since each Newton step reduces to approximate a system of two linear parabolic PDEs. Both Newton–SL and Newton–FD follow a “iteration then discretization” strategy: the Newton direction is computed after discretization, and therefore the iterations do not follow the exact gradient of the continuous problem. As a consequence, the ideal quadratic convergence predicted at the continuous level is not fully observed in practice. In contrast, a “discretization then iteration” strategy, such as FD–Newton, builds the Newton iterations directly from the discrete problem and, in our tests, tends to follow more closely the true discrete gradient, sometimes displaying a behaviour closer to quadratic convergence. Nevertheless, one should not expect the “discretization then iteration” strategy to systematically yield a globally smaller numerical error. Discretizing before or after the Newton step introduces different truncation errors, which may counterbalance any local improvement in the convergence rate. Finally, as expected, the Newton–SL scheme exhibits strong robustness in advection-dominated regimes, whereas both Newton–FD and FD–Newton require a stabilization strategy to prevent breakdown. In this regard, the use of a globalized Newton method proved particularly effective, especially for problems with small viscosity coefficients.
Acknowledgments
The first author were partially supported by Italian Ministry of Instruction, University and Research (MIUR) (PRIN Project2022238YY5, “Optimal control problems: analysis, approximation”) and by INdAM–GNCS Project, codice CUP_\_E53C24001950001.